OpenStudy (firejay5):

If f(x) = 3x -5, then what's f(a)

mathslover (mathslover):

put x=a

OpenStudy (firejay5):

is a 1

mathslover (mathslover):

that is: f(a) = 3(a)-5

OpenStudy (firejay5):

what would it be???

mathslover (mathslover):

What are you asking for? Can you explain exactly?

OpenStudy (firejay5):

you are replacing x and replacing in it for x

OpenStudy (firejay5):

A I meant

mathslover (mathslover):

I replaced X with a

OpenStudy (firejay5):

How would you show your work

mathslover (mathslover):

For example: given f(x) =3(x)^2 Find f(a)

mathslover (mathslover):

first of all replace X by a that is: f(a) = 3(a)^2

mathslover (mathslover):

Just like this

OpenStudy (firejay5):

is A like 1A

OpenStudy (firejay5):

uh

OpenStudy (anonymous):

"a" is just a different variable instead of "x". It's just a name... mostly, functions use 'x' but that's just a habit/tradition. The question wants you to see that it doesn't really matter... that a function of "a" looks just like the function of "x" -- you just replace all the x terms with an "a" in place of the "x".

OpenStudy (anonymous):

So if f(x) = 2x, then f(a) = 2a, and f(y) = 2y and so on. The function "f" says, "take what is in the parenthesis and multiply it by 2"

OpenStudy (firejay5):

so what would it be

OpenStudy (anonymous):

If f(x) = 3x - 5 then f(a) = 3a - 5 In this case, the function "f" just says , "take the thing in the parenthesis, multiply by 3, then subtract 5 from the answer" It doesn't matter if you call the "thing in parenthesis" an "x" or an "a"